OpenStudy (anonymous):

What are the square roots of 196? I know one of them is 14

OpenStudy (accessdenied):

A note is that \( \sqrt{196} \) denotes only a principle square root, while when you solve this equation: \( y^2 = 196 \), (the definition of the square root), you can find two values here by solving...

OpenStudy (anonymous):

Oh so thats where I went wrong

OpenStudy (mrnood):

just a hint: what is \[-1^{2}\]

OpenStudy (anonymous):

Um I can't remember at the moment

OpenStudy (mrnood):

When you remember it will help you with your original question y^2 = 196

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

if your done with your question... please "close" it

OpenStudy (anonymous):

I know I am not done yet I doing the math and then gonna let someone check it

OpenStudy (anonymous):

Ok then, sorry

OpenStudy (anonymous):

just passing some reminders

OpenStudy (anonymous):

I know, and thanks

OpenStudy (mrnood):

You appear to have 'done most of the maths. You need to address the question I asked just now.

OpenStudy (anonymous):

I can't seem to get it.

OpenStudy (accessdenied):

might want to write (-1)^2 in parentheses to be more clear

OpenStudy (anonymous):

I've tried over and over

OpenStudy (mrnood):

\[(-1)^{2}= ?\] \[(1)^{2}= ?\]

OpenStudy (anonymous):

Okay thanks

OpenStudy (mrnood):

Have you got it - or have you given up?

OpenStudy (anonymous):

I got -14

OpenStudy (mrnood):

Well done. So the TWO roots are??

OpenStudy (anonymous):

14 and -14

OpenStudy (mrnood):

Excellent - remember there is always a negative square root of the same value as the positive.

OpenStudy (anonymous):

Okay thanks!